simplify by combining like terms
sqrt 63 - 2 sqrt 28 + 5 sqrt 7

Answer by venugopalramana(3286)

simplify by combining like terms
sqrt 63 - 2 sqrt 28 + 5 sqrt 7=sqrt(9*7)-2sqrt(4*7)+5sqrt(7)
=sqrt(9)sqrt(7)-2sqrt(4)sqrt(7)+5sqrt(7)=3sqrt(7)-2*2sqrt(7)+5sqrt(7)
=sqrt(7){3-4+5}=4sqrt(7)